OpenStudy (anonymous):

Find the equation of the tangent line of the given function at the given point. g(x) =1/square root x , at (1,1)

OpenStudy (ribhu):

\[g(x)= 1/\sqrt{x}\] \[g'(x)= -1/2\sqrt{x}\] slope of tangent at 1,1 will be -1/2 so the equation of tangent becomes (y-1)= -1/2 (x-1) x+2y=3 (final equation)